Solve the system of equations below using elimination by addition.
What are the solutions?
-2x+2y=1
18x-18y=-8

Answer by Boreal(15235)

multiply the top by 9
-18x+18y=9
18x-18y=-8
0=-1
no solution.
Write in slope intercept form
2y=-2x+1 and y =x+(1/2) slope 1
18y=18x+8 and y=x+(4/9) slope 1
both lines have same slope, so they are parallel and don't intersect.